Nicholas1padilla1234:

Solve for x. 12:15 = x:5

Nicholas1padilla1234:

Please help me

KyanTheDoodle:

To be honest, I don't know if there's like a more orthodox way to do this, but this will work too: try to get these ratios to the same greatest common factor.

KyanTheDoodle:

In this case, you would divide the first ratio by 3 so that it's the same with the ratio on the other side.

snowflake0531:

__:__ = __:__ can also b written as fractions \[\frac{12}{15}=\frac{x}{5}\] cross multiply \[12(5) = 15(x)\\60=15x\] then divide 60 by 15